OpenStudy (anonymous):

6. In , a = 6, b = 7, and c = 8. Find the measure of angle A, correct to the nearest degree.

OpenStudy (anonymous):

can someone please explain it to me step by step

OpenStudy (anonymous):

law of cosines to start

OpenStudy (anonymous):

okay! then

OpenStudy (anonymous):

\[a^2=b^2+c^2-2bc\cos(A)\] \[\frac{b^2+c^2-a^2}{2bc}=\cos(A)\]

OpenStudy (anonymous):

once you have \[\cos(A)\] as a number , take the inverse cosine to get the answer we can do it if you like

OpenStudy (anonymous):

i now what the law of cos are soo then all i do is pulg it in

OpenStudy (anonymous):

in other words the answer is \[A=\cos^{-1}(\frac{b^2+c^2-a^2}{2bc})\]

OpenStudy (anonymous):

yup plug it in

OpenStudy (anonymous):

kk thankxx

OpenStudy (anonymous):

looks like about 46.57 degrees http://www.wolframalpha.com/input/?i=arcos%28%287^2%2B8^2-6^2%29%2F%282*8*7%29%29

OpenStudy (anonymous):

lol thankx